OpenStudy (anonymous):

What is the equation of the line that contains the points (3, –6) and (–3, 0)?

OpenStudy (anonymous):

@camerondoherty can you help me

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

\[m=\frac{ 3+3 }{ -6-0 }=\frac{ 6 }{ -6 }=-1\] \[y-y1=m(x-x1)\] y-0=-1(x+3) y=-x-3